OpenStudy (anonymous):

show analytically that f(x)=0 is odd and even

OpenStudy (turingtest):

what is f(-x) ?

OpenStudy (anonymous):

that is why i am here not sure my self but know from reading that this is one of the only functions that is both even and odd, but i have no idea how to show it.

OpenStudy (turingtest):

no matter what x you plug in you get the same number back from the function, right? it's always the same if it is f(x)=0 (or any other constant for that matter)

OpenStudy (turingtest):

if f(x)=0 then f(1)=f(315)=f(-2)=0 yes?

OpenStudy (anonymous):

as i understand it f(1)would be the value to replace x with in an equation and so on for f(315) f(-2) correct?

OpenStudy (turingtest):

right, but in the function f(x)=0 there is no x on the right hand side to replace, so whatever x you plug in changes nothing understand?

OpenStudy (anonymous):

f(x)=0 f(-x)=0(-whatever)=0?

OpenStudy (turingtest):

well yeah... f(x)=0 and f(-x)=0 hence f(x)=f(-x) which means the function is...?

OpenStudy (anonymous):

both even and odd

OpenStudy (turingtest):

well we have only technically shown one, which one?

OpenStudy (anonymous):

even

OpenStudy (turingtest):

correct, because we showed that f(x)=f(-x) so now go ahead and try to prove the oddness yourself

OpenStudy (anonymous):

-f(x)=0 =(-1)0(whatever)=0?

OpenStudy (turingtest):

not quite, but close the (whatever) throws me off

OpenStudy (turingtest):

what is the main thing we need to show for a function to be odd?

OpenStudy (anonymous):

f(x) multiplied by a (-1)

OpenStudy (turingtest):

no we need to show that f(-x)=-f(x)

OpenStudy (turingtest):

so find f(-x) and -f(x) then prove they are equal

OpenStudy (anonymous):

we found f(-x)=0 -f(x)=0(-(x)) if we had a x.

OpenStudy (turingtest):

how did we know that f(-x)=0 ?

OpenStudy (anonymous):

f(x)=0h for the purpose of this disscussion f(-x)=0(-h) f(-x)=0 -f(x)=0(-(h)) even since (-(h) is oppisite of our (h)

OpenStudy (anonymous):

odd

OpenStudy (turingtest):

writing out the stuff (h) in parentheses is making this way confusing...

OpenStudy (turingtest):

we know that f(-x)=0 because if we "plug in" ANY number for x into the function we still get 0 hence the negative of any number also gives 0 since f(x)=0, this means that f(-x) also is zero

OpenStudy (turingtest):

(yes, I know there is nothing to "plug in" in the equation f(x)=0 but try not to get to worried about that, just use the definitions of even and odd)

OpenStudy (turingtest):

since we showed above that f(x)=0 and f(-x)=0 then f(x)=f(-x) now, we have established that the value of the function is \(independent\) of x...

OpenStudy (anonymous):

f(x)=0 f(-x)=-0=0 -f(x)-(0)=0 better

OpenStudy (turingtest):

better, but you are an equals sign short of a perfect proof

OpenStudy (turingtest):

f(x)=0 f(-x)=0 (no -0 because there is no place to "plug in -x, so don't even use it!) -f(x)=-(0)=0

OpenStudy (anonymous):

well that explains it then have even got to proof yet but would welcome any help!

OpenStudy (turingtest):

we need to show two things: that f(-x)=f(x) and that f(-x)=-f(x) the first we can show by "plugging in x" as I described above thing 1: f(x)=0 f(-x)=0 therefor f(-x)=f(x) (even)

OpenStudy (turingtest):

thing 2 f(-x)=0 -f(x)=-(0)=0 but then we observe that this is the same as f(-x), therefor f(-x)=-f(x) (odd)

OpenStudy (anonymous):

ok i see it thank you now just need to get this down mentally, thanks again.

OpenStudy (turingtest):

welcome :)